OpenStudy (anonymous):

could someone explain how to slove this equation im really stuck Cosθ tanθ please will give medal

OpenStudy (anonymous):

use the identity \(\tan( \theta )=\dfrac{\sin(\theta)}{\cos(\theta)}\)

OpenStudy (anonymous):

ok